OpenStudy (arwym):

I am not sure I understand the following instructions on my book (Calculus: The Classic Edition, by Swokowski): "[a] Use Definition (3.1) to find the slope of the tangent line to the graph of the equation at the point of x-coordinate a. [b] Find an equation of the tangent line at P. [c] Sketch the graph and the tangent line at P." Then, I am given: y = square root of x; P(4, 2) The Definition (3.1) refers to the following: limit when h approaches 0 of [ f(a + h) - f(a) ] / h The answer book says that [a] is 1/(2*sqrt(a)) and [b] is y = (1/4)x + 1

OpenStudy (arwym):

What I need is to be sure I understand the instructions correctly. I was trying to use the function y = sqrt(x) in the Definition (3.1), like this: limit when h approaches 0 of [(sqrt(a + h)) - sqrt(a)] / h Am I correct?

OpenStudy (anonymous):

yes that is what you want, although in this case you can do it with numbers

OpenStudy (anonymous):

the point is \[(4,2)\]so you can compute \[\lim_{h\to 0}\frac{\sqrt{4+h}-2}{h}\]

OpenStudy (anonymous):

usual gimmick is to multiply by the conjugate on the top and bottom. you got it from here ?

OpenStudy (arwym):

 This is where I get confused.